An arrow is shot from a bow straight upwards at an initial speed of 104 m/s. How high in the air does the arrow fly before it comes back down?
1
Expert's answer
2020-11-27T10:06:15-0500
"v^2=2gh\\104^2=2(9.8)h\\h=552\ m"